what volume of 3mol h2so4 is required to react with 10g of zinc


1
Expert's answer
2021-05-31T02:38:50-0400

H2SO4 + Zn = ZnSO4 + H2

M (Zn) = 65.4 g/mol

n = m / M

n (Zn) = 10 / 65.4 = 0.15 mol

Acording to the equation, n (Zn) = n (H2SO4) = 0.15 mol

CM = n / V

V (H2SO4) = n / CM = 0.15 / 3 = 0.05 L